# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-edge AI Applications

### v1

@@ Abstract
Modern cutting-edge AI applications are being developed over fast-evolving, heterogeneous, nascent hardware devices. This requires frequent reworking of the AI software stack to adopt bottom-up changes from new hardware, which takes time for general-purpose software libraries. Consequently, real applications often develop custom software stacks optimized for their specific workloads and hardware. Custom stacks help quick development and optimization, but incur a lot of redundant efforts across applications in writing non-portable code. This paper discusses an alternative communication library interface for AI applications that offers both portability and performance by reducing redundant efforts while maintaining flexibility for customization. We present MSCCL++, a novel abstraction of GPU communication based on separation of concerns: (1) a primitive interface provides a minimal hardware abstraction as a common ground for software and hardware developers to write custom communication, and (2) higher-level portable interfaces and specialized implementations enable optimization for different hardware environments. This approach makes the primitive interface reusable across applications while enabling highly flexible optimization. Compared to state-of-the-art baselines (NCCL, RCCL, and MSCCL), MSCCL++ achieves speedups of up to 3.8$\times$ for collective communication and up to 15% for real-world AI inference workloads. MSCCL++ is in production of multiple AI services provided by Microsoft Azure, and is also adopted by RCCL, the GPU collective communication library maintained by AMD. MSCCL++ is open-source and available at https://github.com/microsoft/mscclpp.
Now at xAI. Now at Microsoft AI.
1 Introduction
@@
GPU communication has become a key area of optimization for high-performance AI applications. For instance, modern Large Language Models (LLMs) [11, 5] are often designed to run a single inference task across tens of GPUs (spanning a few physical nodes) to distribute computation and minimize the end-to-end latency. Training tasks often scale much further up to tens of thousands of GPUs [11]. GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink [24], xGMI [3], etc.) or inter-node (Ethernet or InfiniBand (IB) [1]) links. Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distributes and gathers data across all GPUs simultaneously. For instance, communication patterns like AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter [38] are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
@@
A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads. For example, Mixture-of-Experts [14, 12] layers with experts distributed across 16 GPUs over two nodes may spend 40% of end-to-end training time for AllToAll communication. Similarly, a GPT-3 [5] model inference may spend 30% of the end-to-end time for AllReduce communication. A vast amount of efforts has been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) [32, 7, 18], overlapping communication with computation [15, 37], and other lower-level optimizations in the stack [21, 4].
@@
Despite these efforts, achieving high communication performance is still challenging and time-consuming in real applications. Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch. For instance, TensorRT-LLM [25], a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch. This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) [21]. The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes. This raises a simple question: why should it be a new ad-hoc stack from scratch, instead of developing on top of an existing stack?
@@
Surprisingly, such ad-hoc software development has already become a widely common practice in cutting-edge AI applications, not only in GPU communication but in general AI workloads. From our empirical studies, we find that the root cause lies in the fast evolving hardware. The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, which comes up with powerful but immature hardware that is dramatically different from the previous generation. Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency. Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform only sub-optimal for those applications, because it takes time for those libraries to optimize the wide range of workload scenarios over diverse hardware environments. Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments. This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
In this paper, we discuss alternative interface design for AI software libraries to support such custom optimization practices by reducing redundant efforts without harming the flexibility of customization. The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions. Specifically, we have two key observations on what software developers need in real practices. First, software developers usually work closely with hardware experts to design optimization strategies based on the latest hardware features. Therefore, we need a low-level interface that minimally abstracts the hardware, which provides a common ground where both software and hardware developers easily understand and co-work on. Second, most of the existing abstractions and optimizations made in the software stack are not useful or even hinder optimization in real practices. As developers spend most of their time on new hardware and workloads, we need a highly flexible interface where they can try various design and optimization strategies over, rather than sophisticated abstractions based on existing hardware and workloads.
@@
Based on this insight, we propose MSCCL++, [^3] Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus. a novel GPU communication stack designed for high-performance AI applications. Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API). This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, hence quick adoption of latest hardware features with low efforts. The primitive interface abstracts straightforward concepts of communication such as put, get, signal, and wait. This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication algorithms. The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks. It provides zero-copy, one-sided, and asynchronous communication abstractions for efficient communication both within and across nodes. These abstractions enable communication to be efficiently fused within compute kernels, allowing fine-grained overlapping between compute and communication.
@@
While the primitive interface offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces over the primitive interface for quick and easy optimization. Inspired by MSCCLang [10], we design the MSCCL++ stack to separate the declaration of communication algorithms (MSCCL++ DSL (domain-specific language)) from the underlying implementation (MSCCL++ DSL Executor) using an easy-to-use interface (MSCCL++ DSL API). The MSCCL++ DSL accelerates developing custom communication algorithms in new hardware environments, while the performance is still on par with hand-written implementations in most cases. Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is over the MSCCL++ stack (MSCCL++ Collective API) so that applications can adopt it without changing the code at all. Although the power of MSCCL++ is fully realized with application-specific optimization that cannot be delivered with this API, we still observe significant performance benefits even for general collective communication operations.
2 Background and Motivation
2.1 Collective Communication for AI
@@
Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner. They distribute the model and data across multiple GPUs to reduce the computation and memory requirement per GPU. This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations. AllReduce is a common collective communication operation that sums up the partial results from all GPUs and broadcasts the computed result to all GPUs. AllReduce can be divided into two other collective operations: ReduceScatter and AllGather. ReduceScatter sums input buffers on all GPUs and distributes the output buffer equally on all GPUs. AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU. ReduceScatter and AllGather are often used separately depending on the ML model architectures.
2.2 Limitations of Existing Abstractions
@@
Existing collective communication libraries internally implement conventional networking abstractions such as send and recv inside GPU kernels. This would be a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner. However, this would not be the best choice for performance, because such software abstractions are not well aligned with the GPU hardware abstractions. Instead, this paper proposes an API design from the perspective of GPU programming, which makes it more straightforward to efficiently utilize the GPU hardware.
To be specific, we provide an overview of the NCCL architecture in Section 2.2.1 and discuss its inefficiencies in Section 2.2.2. This would be a representative example of existing collective communication libraries for GPUs, because other popular libraries, such as ROCm Collective Communication Library (RCCL) [4] and Microsoft Collective Communication Library (MSCCL) [9], are designed based on NCCL and share the same limitations.
2.2.1 NCCL Architecture
@@
NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs. All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution. ML frameworks, such as PyTorch [27] and TensorFlow [2], wrap the NCCL API in Python, which are called by an ML model.
@@
Figure 1: Ring ReduceScatter kernel in NCCL.
Figure 2: Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB). Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq). GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
@@
NCCL Initialization. During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like InfiniBand (IB) and Ethernet. NCCL uses this information to create communication topologies such as ring and tree with bi-directional links. Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU. The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
@@
NCCL Kernel. NCCL implements a GPU kernel for all collective operations and communication algorithms (ring, tree, etc.). It chooses the best performing algorithm based on the data size of collective communication, namely the message size. NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce. (i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send primitive has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer. Subsequent calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer. Additionally, NCCL defines extra operations that performs multiple of primitives in a single function call but that does not improve over the limitations we discuss in Section 2.2.2.
@@
As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives. [^4] The kernel has been simplified for readability. The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer. The kernel obtains pointers to send and receive buffers, size of the buffer, and initializes a primitive object with information about both previous and next GPUs in the ring (lines 25). The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724). The loop first copies each batch of input data and send it to the next GPU (lines 911). Then the loop goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419). Finally, the GPU stores the data for offset on this GPU (line 24).
2.2.2 Limitations of NCCL Primitives
@@
The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently. As GPU threads are executed in groups (warps in CUDA), [^5] Unlike CPU where each thread runs on a separate context and can be context-switched, GPU threads are grouped and executed altogether for a single instruction. Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth. GPU kernels usually case-by-case optimize the workload to carefully distribute tasks among threads and minimize synchronization overhead. Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call. This static abstraction is not flexible enough to optimize the workload for different links and data sizes, and may introduce unnecessary overheads. We discuss the limitations of NCCL primitives below.
@@
Interconnect Optimizations. A link can allow different ways of transferring data. However, NCCL supports only one mode of data transfer for each link. For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice. Comparing to that of DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth. For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%). However, NCCL uses only thread-copy unless it is infeasible by hardware. [^6] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration. Moreover, using DMA-copy frees GPU threads to other work, such as computations, which can lead to better overlapping of computation and communication depending on the application.
@@
Inflexible Synchronization. NCCL primitives are self-synchronized to ensure data consistency. However, such synchronizations are often too conservative to realize correct communication semantics. Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and start all over again. When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer. To reduce synchronization overhead, we can use two buffers in rotation and only keep first the type of barriers which still ensures consistency and prevents overwriting, but is not possible with self-synchronized NCCL primitives.
Co-optimization with Computation. The most common practices in ML frameworks call separate GPU kernels for computation (e.g., cuBLAS or cuDNN) and collective communication (e.g., NCCL). However, many recent works [29, 15, 12, 17, 8, 39, 37] have shown performance gains by breaking the barrier and co-optimizing computation and communication operations. Unfortunately, coupling NCCL kernels with computation kernels is challenging because NCCL does not design the send and recv primitives as a programmable interface. Since the primitives are blocking and use a FIFO queueing mechanism, making them directly work with other parts of the application is not straightforward and necessitates a memory barrier in between. A lower-level GPU communication abstraction would be needed to enable the co-optimizations.
2.3 Limitations of Existing User Interfaces

### v2

@@ Abstract.
Modern cutting-edge AI applications are being developed over fast-evolving, heterogeneous, nascent hardware devices. This requires frequent reworking of the AI software stack to adopt bottom-up changes from new hardware, which takes time for general-purpose software libraries. Consequently, real applications often develop custom software stacks optimized for their specific workloads and hardware. Custom stacks help in quick development and optimization, but incur a lot of redundant efforts across applications in writing non-portable code. This paper discusses an alternative communication library interface for AI applications that offers both portability and performance by reducing redundant efforts while maintaining flexibility for customization. We present MSCCL++, a novel abstraction of GPU communication based on separation of concerns: (1) a primitive interface provides a minimal hardware abstraction as a common ground for software and hardware developers to write custom communication, and (2) higher-level portable interfaces and specialized implementations enable optimization for different workloads and hardware environments. This approach makes the primitive interface reusable across applications while enabling highly flexible optimization. Compared to state-of-the-art baselines (NCCL, RCCL, and MSCCL), MSCCL++ achieves speedups of up to 5.4$\times$ for collective communication and up to 15% for real-world AI inference workloads. MSCCL++ is in production of multiple AI services provided by Microsoft Azure, and is also adopted by RCCL, the GPU collective communication library maintained by AMD. MSCCL++ is open-source and available at https://github.com/microsoft/mscclpp.
conference: ; ; Now at Microsoft AI. Now at xAI.
1. Introduction
@@
GPU communication has become a key area of optimization for high-performance AI applications. For instance, modern Large Language Models (LLMs) (Dubey et al., 2024; Brown et al., 2020) are often designed to run a single inference task on tens of GPUs (spanning a few physical nodes) to distribute computation and minimize end-to-end latency. Training tasks often scale much further up to tens of thousands of GPUs (Dubey et al., 2024). GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink (NVIDIA, 2025d), xGMI (AMD, 2023), etc.) or inter-node (Ethernet or InfiniBand (IB) (inf, 2023)) links. Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distribute and gather data across all GPUs simultaneously. For example, communication patterns such as AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter (Wikipedia, 2025) are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
@@
A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads. For example, a GPT-3 (Brown et al., 2020) model inference may spend 30% of the end-to-end time in AllReduce communication. Many efforts have been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025a; AMD, 2025).
@@
Despite these efforts, achieving high communication performance is still challenging and time consuming in real applications. Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch. For instance, TensorRT-LLM (NVIDIA, 2025e), a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch. This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025a). The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes. This raises a simple question: Why build a new ad hoc stack from scratch, instead of developing on top of an existing stack?
@@
Surprisingly, such ad hoc software development has already become a widely common practice in cutting-edge AI applications. From our empirical studies, we find that the root cause lies in the fast evolving hardware. The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, resulting in powerful but immature hardware that is dramatically different from the previous generation. Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency. Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform sub-optimally for those applications because it takes time for those libraries to optimize for diverse workload scenarios. Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments. This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
In this paper, we discuss an alternative interface design for AI software libraries to support portability and custom optimizations by reducing redundant efforts without harming the flexibility of customization. The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions based on two observations. First, existing hardware abstractions like NCCL primitives try to hide hardware features to make it easier to program, but may hinder optimizations. Instead, a low-level, primitive interface that minimally abstracts the hardware would make it easier for: (1) application developers to optimize and fine-tune by enabling precise control of the hardware, and (2) library developers to support new hardware features by only providing a shallow layer of abstraction over it. Second, existing collective communication libraries (like NCCL) and DSLs (domain-specific languages; like MSCCLang (Cowan et al., 2023)) provide ease of use, but may lack workload-specific optimizations. Such higher level abstractions could provide portability and enable users to trade-off performance for programming effort.
@@
Based on these insights, we propose MSCCL++ (Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus), a novel GPU communication stack designed for high-performance AI applications. Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API). This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, thus enabling quick adoption of latest hardware features with low effort. The primitive interface abstracts straightforward concepts of communication such as put, signal, wait, and flush. This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication. The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks. It provides zero-copy, one-sided, and asynchronous abstractions for efficient communication both within and across nodes.
@@
While the MSCCL++ Primitive API offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces that are easier to program. Inspired by MSCCLang (Cowan et al., 2023), we separate the declaration of communication algorithms, done using the MSCCL++ DSL API, from the underlying implementation of the algorithm, done by the MSCCL++ DSL Executor using MSCCL++ primitives. The MSCCL++ DSL accelerates the development of application-specific communication algorithms in new hardware environments. Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is (MSCCL++ Collective API) using MSCCL++ primitives so that applications can adopt it without changing the code at all.
In our evaluation, we compare SpeCL with NCCL, ROCm Collective Communication Library (RCCL) (AMD, 2025) and Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) on NVIDIA A100, NVIDIA H100, and AMD MI300x GPUs. We show that collective communication kernels like AllReduce and AllGather written using the MSCCL++ DSL API are up to 5.4×5.4$\times$5.4 × faster than the other libraries. We also re-implement some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their counterparts. These algorithm implementations are included in the MSCCL++ Collective library. These benefits translate to AI inference workloads, yielding up to $15\%$ speedup in decodes over NCCL for a popular LLM, Llama2-70b (Touvron et al., 2023).
2. Background and Motivation
2.1. Collective Communication for AI
@@
Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner. They distribute the model and data across multiple GPUs to reduce the computation and memory requirement per GPU. This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations. AllReduce is a common collective communication operation that sums the partial results of all GPUs and broadcasts the computed result to all GPUs. AllReduce can be divided into two other collective operations: ReduceScatter and AllGather. ReduceScatter sums the input buffers on all GPUs and distributes the output buffer equally on all GPUs. AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU. ReduceScatter and AllGather are often used separately depending on the ML model architectures.
2.2. Existing Hardware Abstractions
@@
Existing collective communication libraries implement conventional networking abstractions such as send and recv inside GPU kernels. This is a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner. However, it is not the best choice for performance because such software abstractions are not well aligned with the GPU hardware abstractions. Instead, this paper proposes an API design from the perspective of GPU programming, which makes it easier to efficiently utilize GPU hardware. In this section, we provide an overview of the NCCL architecture and discuss its inefficiencies. This is representative of existing collective communication libraries for GPUs, because other popular libraries, such as RCCL (AMD, 2025) and MSCCL (Cowan et al., 2022), are designed based on NCCL and share the same limitations.
2.2.1. NCCL Architecture
@@
NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs. All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution. ML frameworks, such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016), wrap the NCCL API in Python, which are called by an ML model.
@@
Figure 1. Ring ReduceScatter (simplified) in NCCL.
@@
NCCL Initialization. During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like IB and Ethernet. NCCL uses this information to create communication topologies such as ring and tree with bi-directional links. Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU. Instead, users can also register their own send and receive buffers (using ncclMemAlloc()). The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
@@
NCCL Kernel. NCCL implements a GPU kernel for all collective operations and communication algorithms (ring, tree, etc.). It chooses the best performing algorithm based on the data size of collective communication, namely the message size. NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce. (i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer. Calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer. Additionally, NCCL defines extra operations that execute multiple primitives in a single fused function. However, this does not improve over the limitations we discuss in Section 2.2.2.
@@
As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives. The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer. The kernel obtains pointers to send and receive buffers, size of the buffer, and initializes a primitive object with information about both previous and next GPUs in the ring (lines 25). The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724). The loop first copies each batch of input data and sends it to the next GPU (lines 911). The loop then goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores it in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419). Finally, the GPU stores the data for the offset on this GPU (line 24).
2.2.2. Limitations of NCCL Primitives
@@
The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently. As GPU threads are executed in groups (warps in CUDA), [^3] Unlike CPU where each thread runs on a separate context and can be context-switched, GPU threads are grouped and executed altogether for a single instruction. Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth. GPU kernels usually optimize the workload case-by-case to carefully distribute tasks among threads and minimize synchronization overhead. Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call. This static abstraction is not flexible enough to optimize the workload for different links and data sizes, and may introduce unnecessary overheads. We discuss the limitations of NCCL primitives below.
@@
Inflexible Synchronization. To ensure data consistency, the NCCL primitives are designed to be self-synchronous. Such synchronizations are often too conservative to realize correct communication semantics. Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and starts all over again. When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer. To reduce synchronization overhead, we can use two buffers in rotation and only keep the first type of barriers which is sufficient to prevent overwriting. However, this is not possible with self-synchronized NCCL primitives.
Figure 2. Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB). Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq). GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
Interconnect Optimizations. A link can allow different ways of transferring data. However, NCCL supports only one mode of data transfer for each link. For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice. Compared to DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth. For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%). However, NCCL uses only thread-copy unless it is infeasible by hardware. [^4] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration. Moreover, using DMA-copy frees GPU threads to do other work.
2.3. Existing Communication Libraries